### v1

# Leveraging Modality Tags for Enhanced Cross-Modal Video Retrieval

Abstract
Video retrieval requires aligning visual content with corresponding natural language descriptions. In this paper, we introduce Modality Auxiliary Concepts for Video Retrieval (MAC-VR), a novel approach that leverages modality-specific tags – automatically extracted from foundation models – to enhance video retrieval. We propose to align modalities in a latent space, along with learning and aligning auxiliary latent concepts, derived from the features of a video and its corresponding caption. We introduce these auxiliary concepts to improve the alignment of visual and textual latent concepts, and so are able to distinguish concepts from one other.
We conduct extensive experiments on five diverse datasets: MSR-VTT, DiDeMo, TGIF, Charades and YouCook2. The experimental results consistently demonstrate that modality-specific tags improve cross-modal alignment, outperforming current state-of-the-art methods across three datasets and performing comparably or better across the other two.
I Introduction
The emergence of prominent video-sharing platforms like YouTube and TikTok has supported uploading of millions of videos daily. The demand for better video retrieval methods, which align textual queries with relevant video content, has subsequently increased. Most existing works use two main approaches. The first [1, 2, 3] exclusively uses word and frame features. However, the second [4, 5, 6, 7, 8, 9] uses additional multi-modal information from videos – such as audio, speech, and objects – that is encoded and used for feature aggregation. In real-world scenarios, online videos often come with related textual information, such as tags – keywords associated with a video that describe its content and make it easier to search/filter. Few works [10, 11, 12] extract and exploit tags in video retrieval to better align the visual and textual modalities. These works only focus on extracting tags from the visual modality using pre-trained experts or a predefined visual token vocabulary. Inspired by these, we develop a novel method called MAC-VR that integrates multi-modal information by independently extracting relevant tags for both videos and texts without any manual annotation, utilizing the extensive knowledge from pre-trained Vision-Language Models (VLM) and Large Language Models (LLM) as shown in Fig. 1. The example in Fig. 1 shows the query “a girl doing gymnastics in the front yard”, the extracted visual (VT) and textual (TT) tags include sports, physical, outdoors, and outside can help align this video to the corresponding caption.
Figure 1: Tags are extracted from both videos by VLM and texts by LLM using custom prompts designed to generate the most relevant tags for each modality. For example, visual and textual tags like sports and physical can help align this video to the corresponding caption. We learn latent auxiliary concepts from these tags, that help align the videos and texts.
We build on the recent work DiCoSA [3], due to its performance and efficiency, where the visual and textual coarse features are split into compact latent factors which explicitly encode visual and textual concepts. Our MAC-VR extends this work by introducing additional (i.e. auxiliary) modality-specific latent concepts. We learn these auxiliary concepts from visual and textual tags using modality-specific foundation models. These are aligned to the latent concepts directly extracted from video and text, through an introduced Alignment Loss.
Recently, many works [13, 3, 14, 15, 16, 10] use different inference strategies to improve the final video retrieval performance such as Querybank Normalisation (QB) [17] and Dual Softmax (DSL) [18]. In this paper, we analyse the impact of such strategies on our MAC-VR architecture to ensure fair comparison with state-of-the-art (SOTA) methods. Our results show that auxiliary concepts of both modalities, in addition to the Alignment Loss, help boost the retrieval performance and better distinguish the latent concepts.
Our contribution is as follows: (i) We propose automatic extraction of modality-specific tags from foundational models to augment the video/text modalities. (ii) We use the tags to learn auxiliary latent concepts in each modality to extract meaningful representations. (iii) We propose a new Alignment Loss to better align and distinguish these learnt latent concepts. (iv) We analyse the impact of different inference strategies, extensively and fairly comparing our proposal with SOTA methods. (v) We conduct experiments on five datasets: MSR-VTT, DiDeMo, TGIF, Charades, and YouCook2. Across all datasets, the addition of our auxiliary concepts improves performance over our baseline. A detailed ablation on MSR-VTT verifies our design.
Figure 2: Examples of visual and textual tags (middle) for videos (left) and corresponding captions (right). Blue and orange highlight meaningful tags.
II Related works
Video Retrieval. Video retrieval learns an embedding for video and text to establish effective connections between video content and natural language descriptions. Early approaches [4, 5, 6, 7, 8, 9, 19, 20, 21, 22] relied on pre-trained features and/or multi-modal information inherent in videos, such as audio or speech, specialized to bridge the gap between videos and text. Notably, MMT [5] explores multi-modal data extracted by seven pre-trained experts but integrates them without explicit guidance, employing a brute-force method. Input modalities have also been masked, e.g. in [7], where the method can learn robust representations that enhance cross-modal matching; or video embeddings, such as in [19], by local temporal context, i.e. neighbouring actions. On the contrary, MAC-VR uses only video and text modalities without considering any additional modalities, such as audio or speech, or context.
Recent advancements in video retrieval have followed two main methodologies. The first involves extensive pre-training of models on large-scale video-text datasets, [23, 24]. The second focuses on transferring knowledge from image-based CLIP models [25] trained on extensive image-text pairs [21, 1, 26, 2, 14, 27, 28, 29, 30, 31, 3, 32, 33]. Some works [28, 30] use a distillation approach where a large network is first trained as a teacher network and then a smaller network is trained as a student network. In contrast, MAC-VR does not use any distillation approach and is trained directly by introducing auxiliary modality-specific tags. Similar to [3], where learnable queries and latent concepts are learnt during training, we learn latent auxiliary concepts from our modality-specific tags in addition to visual and textual latent concepts and use them as additional features to align the visual and textual concepts.
Vision-Language and Large Language Models in Image and Video Retrieval. The integration of Vision-Language Models (VLM) [34, 35, 36, 37, 38] and Large Language Models (LLM) [39, 40, 41, 42] in image [43, 44, 45, 46, 47] and video retrieval [48, 49, 12, 50, 51, 52, 53] has enabled significant advancements, showcasing the impressive understanding capabilities of these models. In [49] the authors demonstrate that LLMs can enhance the understanding via generation of video content, transferring their rich semantic knowledge. In [48], the authors explore how additional captions can enhance video retrieval by providing richer semantic context and improving matching accuracy between textual queries and video content. In contrast to these works, we do not generate additional captions as in [49, 48] but we leverage pre-trained VLMs and LLMs to generate words (i.e. visual and textual tags) that highlight relevant aspects of the action shown in the video and described by the caption. Our ablations verify the effectiveness of tags over additional captions extracted from the same foundation models.
Tags in Image and Video Understanding. The notion of tags in image and video understanding has found application across various tasks, including Video Retrieval [11, 10, 12], Video Moment Retrieval [54, 55], Video Recognition [56, 57], Fashion Image Retrieval [58, 59, 60, 61, 62] and Image Retrieval [63, 64, 65, 66, 67]. In [11, 10, 12], only the visual modality is considered to extract tags. Differently, MAC-VR considers both visual and textual modality to extract tags. More precisely, [11, 10] use pre-trained experts to extract tags from various modalities of videos, including object, person, scene, motion, and audio. In contrast, MAC-VR does not use pre-trained expert models to extract tags from a video or any additional modality such as audio. However, we generate visual and textual tags directly from videos and captions by using VLM and LLM, respectively. Similar to us, [12] uses image-language models to translate the video content into frame captions, objects, attributes, and event phrases by using a pre-defined visual token vocabulary. MAC-VR does not generate any additional captions from frames, and does not have any pre-fixed visual token vocabulary, instead using both the raw video and caption input to extract tags automatically.
III Modality Auxiliary Concepts for Video Retrieval
We first define cross-modal text-to-video retrieval and inference strategies in Sec. III-A before describing our tag extraction approach in Sec. III-B. Finally, in Sec. III-C, we introduce our MAC-VR architecture.
Figure 3: Architecture of MAC-VR: Given a video $v_{i}$ and its corresponding caption $t_{i}$, we generate auxiliary Visual Tags (VT) and textual Tags (TT) using a VLM and an LLM, respectively. A shared text encoder projects the caption and the auxiliary tags $T_{i}$, $A_{i}^{v}$ and $A_{i}^{t}$ to a common space with the Text-Conditioned Video Encoder (T-CVE). Visual $e_{i,k}^{v}$ and textual $e_{i,k}^{t}$ concepts are aligned to each other by the contrastive loss $\mathcal{L}_{C}$ and are aligned to auxiliary visual $e^{a^{v}}_{i,k}$ and textual $e^{a^{t}}_{i,k}$ concepts by our Alignment Loss $\mathcal{L}_{A}$. An MLP then estimates confidence scores for each concept, to compute a weighted sum for the similarity function that is used in our Cross-Modal Loss $\mathcal{L}_{S}$.
III-A Cross-Modal Text-to-Video Retrieval
Given a pair $(v_{i},t_{i})$, where $v_{i}$ represents a video and $t_{i}$ denotes its corresponding caption, the objective of Cross-Modal Text-to-Video Retrieval is to retrieve $v_{i}$ from a large gallery of videos, given the caption $t_{i}$ as query. Typically, models use two projection functions: $f_{v}:v_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ and $f_{t}:t_{i}\longrightarrow\Omega\in\mathbb{R}^{d}$ to map the respective modalities into a shared d-dimensional latent embedding space, $\Omega$. The aim is to align the representations in this space so that the representation of a video is close to that of its corresponding caption. Following training, the standard inference strategy (IS) embeds a test video gallery and ranks these based on their distance from each query caption. Recent approaches utilise additional inference strategies to improve performance: Querybank Normalisation (QB) [17] and Dual Softmax (DSL) [18] are the most popular. We introduce these here and later show their impact across current SOTA.
The QB strategy [17] was introduced to mitigate the hubness problem of high-dimensional embedding spaces [68], where a small subset of samples tends to appear far more frequently among the k-nearest neighbours queries. To mitigate this, the similarities between embeddings are altered to minimise the influence of hubs. A set of samples (the querybank) is sampled from the queries within the training set and used as a probe to measure the hubness of the gallery at test time. More formally, given the vector of unnormalised similarities at test time, $S(v_{j},t_{i})$ and a probe matrix $P$, in which each row is a probe vector of similarities between the querybank and each element in the gallery, we re-weight the query-gallery similarities: $\eta_{q}=\text{QB}(S(v_{j},t_{i}),P)$, where the QB is the querybank normalisation function using Dynamic Inverted Softmax (DIS), introduced in [17].
The second commonly used inference strategy, the DSL strategy [18], introduces an intrinsic prior of each pair in a batch to correct the similarity matrix and achieves the dual optimal match. In practice, we modify the original $S(v_{j},t_{i})$ by multiplying it with a prior $r_{i,j}$. Therefore, we can define the new similarity matrix as $\hat{S}(v_{j},t_{i})=r_{i,j}S(v_{j},t_{i})$, where the prior is defined as $r_{i,j}=\frac{exp(\tau_{r}S(v_{i},t_{i}))}{\sum_{j}exp(\tau_{r}S(v_{i},t_{j}))}$, where $\tau_{r}$ is a temperature hyper-parameter to smooth the gradients. While this strategy can be used both in training and inference, it is now regularly used only during inference.
III-B Tag Extraction
We propose to extract tags automatically from either the video $v_{i}$, using a VLM, or the text in the caption $t_{i}$, using an LLM. These tags are word-level representations of common objects, actions, or general ideas present in the video or the caption. They can add additional useful information to retrieve the correct video given a text query as shown in Fig. 2. For instance, given the caption: a commercial for the Mazda 3 the car sliding around a corner, the general tags estimated from this caption are: product showcase, advertising technique, automotive marketing, which reflect the commercial. These words are abstract terms that go beyond the exact caption but can help the model to understand the specific characteristics of this caption better.
In contrast, leveraging the video modality to create tags enables us to both capture a broader array of visual elements that characterise the video content and also have a representation of the video in words, facilitating matching the video content to the captions within the text modality. E.g., given the video associated with the previous caption, extracted visual tags include road, vehicle, car, transportation, engine, reflecting important objects in the video, and racing, driving reflecting the action in the video. These tags directly correspond to pertinent visual components of the video.
To extract visual and textual tags, we use a custom prompt (see Sec. A in appx.) to query the most relevant general tags for the input video $v_{i}$ and caption $t_{i}$. We extract tags individually from both modalities so they can be used for training and inference. As the tags are extracted from a single modality, which we call modality-specific tags. We detail how these tags are used within MAC-VR next.
III-C Architecture
We start from a standard Text-Conditioned Video Encoder (T-CVE) before incorporating our proposed tags into each modalities’ latent concepts. These concepts are aligned and pooled to find the similarity between a video and a caption. Our proposed architecture is summarised in Fig. 3.
Text-Conditioned Video Encoder (T-CVE)
Given a caption $t_{i}$, we extract its text representation ${T_{i}\in\mathbb{R}^{d}}$. For the video representation, we first sample $N_{v}$ frames from a video $v_{i}$ and then encode them and aggregate the embedding of all frames to obtain the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$. Since captions often describe specific moments, as shown in previous works [69, 3, 26], matching only the relevant frames improves semantic precision and reduces noise. To achieve this, we aggregate the frame representations conditioned on the text. Firstly, we calculate the inner product between the text and the frame representation $F_{j}$ with $j\in\{1,...,N_{v}\}$:
$a_{i,j}=\frac{exp((T_{i})^{\top}F_{j}/\tau_{a})}{\sum^{N_{v}}_{k=1}(exp((T_{i})^{\top}F_{k}/\tau_{a}))}$ (1)
where $\tau_{a}$ is a hyper-parameter that allows control of the textual conditioning. Then, we get the text-conditioned video representation $V_{i}\in R^{d}$ defined as $V_{i}=\sum\limits^{N_{v}}_{k=1}a_{i,k}F_{k}$.
Latent Concepts
To utilise the visual and textual tags extracted from foundational models in Sec III-B, we first randomly pick $N$ visual and textual tags (per modality) during training and order them into two distinct comma-separated sentences that start with “A video of”. We extract visual $A_{i}^{v}$ and textual $A_{i}^{t}$ coarse tag features by using the same text encoder used for the caption. Therefore, given a video/caption pair $(v_{i},t_{i})$ we get a quadruple $(V_{i},T_{i},A_{i}^{v},A_{i}^{t})$. Inspired by [3], we disentangle each element of the quadruple into $K$ independent, equal-sized latent concepts. For example, when disentangling $V_{i}$, we get $K$ independent latent concepts, i.e. $E_{i}^{v}=[e^{v}_{i,1},...,e^{v}_{i,K}]$. Each latent concept $e^{v}_{i,k}\in R^{d/K}$ represents a distinct concept and the independence of these factors ensures that each concept is uncorrelated to the other $K\text{-}1$ latent concepts, and is thus calculated by independently projecting the text representation:
$e^{v}_{i,k}=W^{v}_{k}V_{i}$ (2)
where $W^{v}_{k}$ is a trainable parameter. Similarly, $E_{i}^{t}$, $E_{i}^{a^{v}}$ and $E_{i}^{a^{t}}$ represent the latent concepts of the text $T_{i}$, the visual $A_{i}^{v}$ and the textual $A_{i}^{t}$ tag representations and are calculated in the same way. We name the $K$ latent concepts of the tags representation as the auxiliary visual concepts $E_{i}^{a^{v}}$ and auxiliary textual concepts $E_{i}^{a^{t}}$ respectively. We now have four disentangled representations for visual $E_{i}^{v}$, textual $E_{i}^{t}$, auxiliary visual $E_{i}^{a^{v}}$, and auxiliary textual $E_{i}^{a^{t}}$ concepts. Until now, these subspaces have been disentangled independently. We next describe how the alignment of these latent concepts can be used for enhancing cross-modal retrieval.
#Unique Visual Tags (VT) #Unique Textual Tags (TT)
Datasets #Videos # Captions Train Val Test Train Val Test
MSR-VTT-9k [5] 10,000 190,000 63,383 - 12,118 320,351 - 8,326
MSR-VTT-7k [70] 8,010 141,200 52,309 - 12,118 270,367 - 8,326
DiDeMo [71] 10,642 10,642 50,712 10,924 10,636 34,662 9,234 8,266
TGIF [72] 100,551 100,551 105,895 28,500 29,538 176,353 44,053 45,417
Charades [73] 9,848 9,848 41,538 - 15,360 25,406 - 9,874
YouCook2 [74] 13,829 13,829 33,786 - 16,457 26,035 - 12,227
TABLE I: Statistical analysis of tags after extraction.
Alignment of Disentangled Latent Concepts
By default, approaches such as [3] directly align latent representations of videos and captions through a contrastive loss. Here, we extend this by applying the contrastive loss to align modality-specific concepts to the corresponding auxiliary concepts. In this way, modality-specific concepts assist MAC-VR in better aligning the video and text modalities. Specifically, we consider the visual concepts $E_{i}^{v}$ and the auxiliary visual concepts $E_{i}^{a^{v}}$. For each disentangled concept pair $(e_{i,k}^{v},e_{i,k}^{a^{v}})$ we minimise the distance between this pair then maximise the distance to other disentangled concepts, i.e. $(e_{i,k}^{v},e_{i,l}^{a^{v}});l\neq k$ in a contrastive fashion to align modality concepts. Recall that these latent concepts are learnt, and thus through this alignment, we aim to learn a representation of the video that matches the latent representations of the tags extracted from the VLM.
Similarly, we align the auxiliary textual concepts $E_{i}^{a^{t}}$ to the latent concepts $E_{i}^{t}$ extracted directly from the caption. We combine both modalities’ alignment of latent concepts to auxiliary latent concepts and refer to this as the Alignment Loss $\mathcal{L}_{A}$ which aligns $E_{i}^{v}$ with $E_{i}^{a^{v}}$ and $E_{i}^{t}$ with $E_{i}^{a^{t}}$.
Weighted Similarity and Training Loss
Only a subset of visual concepts are usually described in the corresponding text. Therefore, we cannot directly leverage correlations between their latent concepts, so we use adaptive pooling to define weights for the visual and textual concepts and reduce their impact on the final similarity calculation. To do this, we design an adaptive module to estimate the confidence of each cross-modal concept matching. For each concept $k$, we concatenate the modality and auxiliary modality concepts $[e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}]$ and use them to calculate the confidence of each cross-modal concept matching. We thus calculate:
$c_{i,k}=MLP([e^{v}_{i,k},e^{t}_{i,k},e^{a^{v}}_{i,k},e^{a^{t}}_{i,k}])$ (3)
If the confidence $c_{i,k}$ is low, the corresponding latent concept of the $k^{th}$ subspace is matched with a low score. Given this confidence, we aggregate all the visual and textual latent concept pairs to calculate the overall similarity of the video and text $S(v_{i},t_{i})$, through adaptive pooling as:
$S(v_{i},t_{i})=\sum^{K}_{k=1}c_{i,k}\cdot\frac{(e^{t}_{i,k})^{\top}e^{v}_{i,k}}{|\|e^{t}_{i,k}\|\|e^{v}_{i,k}\|}$ (4)
Following common approaches, we use the InfoNCE loss [75, 76] as our Cross-Modal Loss ($\mathcal{L}_{S}$) to optimise the cross-modal similarity $S(v_{i},t_{i})$; the contrastive loss $\mathcal{L}_{C}$ to align the latent modality concepts (as introduced in [3]); and our proposed Alignment Loss $\mathcal{L}_{A}$ to align the modality with the auxiliary modality concepts:
$\begin{split}\mathcal{L}=\mathcal{L}_{S}(S(v_{i},t_{i}))+\alpha_{1}\mathcal{L}_{C}(E_{i}^{v},E_{i}^{t})+\\ \alpha_{2}\mathcal{L}_{A}(E_{i}^{v},E_{i}^{t},E^{a^{v}}_{i},E^{a^{t}}_{i}).\end{split}$ (5)
where $\alpha_{1}$ and $\alpha_{2}$ are weight parameters. During inference, the weighted similarity $S(v_{i},t_{i})$ as in Eq. 4 is used, given the query caption and every video in the gallery.
Figure 4: Comparison of $RSum$ against baseline MAC-VR. $BS$: Batch Size. $N_{v}$: Number of Frames.
MSR-VTT-9k MSR-VTT-7k DiDeMo
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
CLIP4Clip [77] - Neurocomp.’22 44.5 71.4 15.3 197.5 42.1 71.9 16.2 195.4 43.4 70.2 17.5 194.2
CenterCLIP [78] - SIGIR’22 44.2 71.6 15.1 197.9 43.7 71.3 16.9 195.8 - - - -
X-Pool [26] - CVPR’22 46.9 72.8 14.3 201.9 43.9 72.5 14.6 198.7 - - - -
TS2-Net [13] - ECCV’22 47.0 74.5 13.0 205.3 - - - - 41.8 71.6 14.8 195.4
EMCL-Net [15] - NeurIPS’22 46.8 73.1 - 203.0 - - - - 46.8 74.3 12.3 204.2
EMCL-Net [15] - NeurIPS’22 - - - - 45.2 71.4 15.0 197.8 44.1 71.7 15.1 197.8
VoP [27] - CVPR’23 44.6 69.9 16.3 194.8 42.7 68.2 15.9 190.2 46.4 71.9 13.6 199.8
MuMUR [79] - Inf. Retr. J.’23 46.4 72.6 13.9 201.2 44.8 72.0 - 199.3 44.4 74.3 - 201.8
PiDRo [29] - ICCV’23 48.2 74.9 12.6 206.4 - - - - 48.6 75.9 11.8 208.9
HBI [80] - CVPR’23 48.6 74.6 12.0 206.6 - - - - 46.9 74.9 12.1 204.5
DiffusionRet [14] - ICCV’23 49.0 75.2 12.1 206.9 - - - - 46.7 74.7 14.3 204.1
Prompt Switch [81] - ICCV’23 47.8 73.9 14.4 203.9 - - - - - - - -
Cap4Video [48] - CVPR’23 49.3 74.3 12.0 207.4 - - - - 52.0 79.4 10.5 218.9
UCoFiA [82] - ICCV’23 49.4 72.1 12.9 205.0 - - - - 46.5 74.8 13.4 205.7
UCoFiA [82] - ICCV’23 - - - - 45.2 69.6 15.7 194.2 42.1 69.2 16.3 190.4
PAU [83] - NeurIPS’23 48.5 72.7 14.0 203.7 - - - - 48.6 76.0 12.9 209.1
TABLE [10] - AAAI’23 47.1 74.3 13.4 204.3 - - - - 47.9 74.0 14.3 204.0
UATVR [33] - ICCV’23 47.5 73.9 12.3 204.9 - - - - 43.1 71.8 15.1 197.2
TeachCLIP [30] - CVPR’24 46.8 74.3 - - - - - - 43.7 71.2 - -
MV-Adapter [31] - CVPR’24 46.2 73.2 - 202.1 - - - - 44.3 72.1 - 196.9
BiC-Net [84] - TOMM’24 39.4 75.5 - 201.6 32.8 68.2 - 183.4 - - - -
RAP [85] - ACL’24 44.8 71.4 14.4 197.7 - - - - 42.6 70.4 18.0 192.6
MAC-VR (ours) - - 48.8 74.4 12.3 206.9 45.3 71.9 15.0 199.0 43.4 72.7 16.9 198.4
QB-Norm [17] QB CVPR’22 47.2 73.0 - 203.2 - - - - 43.3 71.4 - 195.5
DiCoSA [3] QB IJCAI’23 47.5 74.7 13.2 206.0 - - - - 45.7 74.6 11.7 203.8
DiffusionRet [14] QB ICCV’23 48.9 75.2 12.1 207.2 - - - - 48.9 75.5 14.1 207.7
MAC-VR (ours) QB - 49.3 75.9 12.3 208.7 45.6 72.7 16.0 200.4 45.5 74.8 16.2 202.6
EMCL-Net [15] DSL NeurIPS’22 51.6 78.1 - 215.0 - - - - - - - -
EMCL-Net [15] DSL NeurIPS’22 - - - - - - - - 47.6 73.5 11.9 203.9
TS2-Net [13] DSL ECCV’22 51.1 76.9 11.7 213.6 - - - - 47.4 74.1 12.9 203.9
TABLE [10] DSL AAAI’23 52.3 78.4 11.4 215.9 - - - - 49.1 75.6 14.8 207.6
UATVR [33] DSL ICCV’23 49.8 76.1 12.3 211.4 - - - - - - - -
RAP [85] DSL ACL’24 - - - - - - - - 47.1 74.1 13.9 203.6
MAC-VR (ours) DSL - 53.2 77.7 10.0 216.2 49.8 74.6 12.7 208.7 50.2 75.2 15.1 209.6
TGIF Charades YouCook2
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
RIVRL [86] - TCSVT’22 12.2 26.8 - 74.4 - - - - - - - -
X-Pool [26] - CVPR’22 - - - - 16.1 35.2 67.2 96.2 - - - -
AME-Net [87] - TOMM’22 - - - - - - - - 7.6 21.5 - 61.9
EMCL-Net [15] - NeurIPS’22 13.2 27.3 327.5 75.3 16.4 36.3 69.5 100.0 11.8 28.8 121.6 80.4
RoME [88] - CoRR’22 - - - - - - - - 6.3 16.9 - 48.4
SwAMP [89] - AISTATS’23 - - - - - - - - 9.4 24.9 - 69.6
MuMUR [79] - Inf. Retr. J.’23 - - - - 16.6 37.5 52.7 104.1 - - - -
UCoFiA [82] - ICCV’23 13.0 27.1 323.8 74.8 16.4 37.2 74.6 101.3 12.4 30.4 123.5 83.8
LTME [90] - ICASSP’24 10.6 24.1 - 66.9 - - - - - - - -
BiC-Net [84] - TOMM’24 - - - - - - - - 8.7 23.9 - 66.1
MAC-VR (ours) - - 14.2 28.8 294.4 79.7 17.8 40.0 55.0 109.2 12.3 30.7 106.7 85.3
EMCL-Net [15] DSL NeurIPS’22 14.8 29.4 317.3 80.8 18.2 38.2 63.1 105.7 12.6 31.7 104.5 86.8
MAC-VR (ours) DSL - 16.2 31.2 284.3 86.4 19.5 42.1 52.4 116.0 13.1 32.1 99.3 88.6
TABLE II: Comparison with SOTA across all the datasets. --: unreported results. our reproduced results. IS: Inference Strategy. The highest $RSum$ in each block is marked in bold.
IV Experiments
First, we introduce datasets and metrics considered to evaluate MAC-VR in Sec.IV-A. Then, we introduce the implementation details for reproducibility in Sec. IV-B. Sec. IV-C and Sec. IV-D show the comparison with the baseline and SOTA works, and ablation experiments respectively.
IV-A Datasets and Metrics
MSR-VTT [91] is commonly utilised for video retrieval. It comprises of 10K YouTube videos, each with $20$ captions. We utilise two commonly used training splits, with the same test set of 1K videos: the 9k-Train split [5] and the 7k-Train split [70]. DiDeMo [71] collects 10K Flickr videos annotated with paragraph captions. This dataset is evaluated in a video-paragraph retrieval manner [2]. TGIF [72] is a large dataset that contains more than 100K animated short GIFs collected from Tumblr, and natural language sentences annotated via crowdsourcing. We use the split proposed in [92]. Charades [73] contains $9,848$ videos of $157$ action classes in total, where each video is associated with a caption. This dataset consists of fine-grained videos of recorded human actions. We use the standard training and test splits [93]. YouCook2 [74] is a fine-grained dataset that contains 14K YouTube cooking videos that cover $89$ recipes. A textual sentence describes each video clip. Following [70], we evaluate this dataset on the validation set.
Modality-Specific Tags. We extract modality-specific tags from all videos and captions in the datasets above. On average we have between $27$ and $29$ visual and textual tags across all the datasets and splits. Tab. I presents the number of unique visual and textual tags across datasets.
Metrics. We present the retrieval performance for text-to-video retrieval task using standard metrics: Recall at $L=1,5$ ($R@L$) and mean rank ($MeanR$). We report the complete tables including $R@10$ and median rank $MR$ in Sec. B in the appx. We use the sum of cross-modal $R@L$ with $L=1,5,10$ as $RSum$ to rank methods.
IV-B Implementation Details
We build our model using DiCoSA [3] as our baseline. Our architecture employs CLIP’s VIT-B/32 [94] as the image encoder and CLIP’s transformer base as the text encoder to encode captions and visual/textual tags. For tag generation, we use the fine-tuned VideoLLaMA2 [37] as the VLM for visual tags and Llama3.1 [95] as the LLM for textual tags. VideoLLaMA2 runs on $8$ sparsely sampled frames with temperature values $\tau\in{0.7,0.8,0.9,1.0}$ – the same values are used for Llama3.1 – to yield diverse tags. During training, we randomly select $N$ tags, while during inference we use fixed $M$ tags for deterministic results; we set $N=M=12$ for training and evaluating MAC-VR, except when using DSL, where $N=6$ and $M=8$.
Following the implementation details of [3], we optimize the model using an Adam optimizer with linear warm-up, setting initial learning rates of 1e-7 for the text and video encoders and 1e-3 for other modules. Unless specified, we set $\tau_{a}=3$ in Eq. 1, $K=8$, and weight parameters $\alpha_{1}=\alpha_{2}=1$. Our MLP consists of two 256-length linear layers with a ReLU activation in between. We use $N_{v}=12$ frames for all datasets except DiDeMo, where $N_{v}=50$ frames are used for video-paragraph retrieval. The model is trained for $10$ epochs with a batch size of $128$ ($64$ for DiDeMo due to GPU memory) and employs a 4-layer transformer to aggregate frame embeddings.
IV-C Results
First, we present the comparison of MAC-VR against the baseline DiCoSA [3]. Then, we compare MAC-VR with SOTA methods, particularly highlighting the impact of inference strategies on fairness of comparison.
Comparison with Baseline. In Fig. 4, we compare MAC-VR with our Baseline DiCoSA trained using the same training parameters, i.e. batch size, $BS$; same number of frames, $N_{v}$; and when using all inference strategies (IS). In all the datasets we improve $\Delta RSum$ compared to our baseline DiCoSA across all IS – by $+0.9$ for MSR-VTT-9k, $+3.4$ for MSR-VTT-7k, $+2.8$ for DiDeMo, $+9.5$ for Charades, $+3.0$ for TGIF and $+0.6$ for YouCook2 when using the DSL approach as inference strategy. The corresponding tables are shown in Sec. B in the appx.
Comparison with SOTA. In Tab. II, we provide a comprehensive comparison of MAC-VR against all SOTA works. To ensure fairness, we divide methods by their inference strategy and consider three different settings: MAC-VR without any inference strategy, MAC-VR with QB, and MAC-VR with DSL.
$R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
baseline 52.1 77.3 12.9 215.3
+VT 52.2 77.3 10.4 215.2
+TT 52.0 77.6 10.4 215.6
+VT+TT 53.2 77.7 10.0 216.2
TABLE III: Ablation on which modality to extract tags from.
Foundation Models
VT TT $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
VL L2 52.0 77.5 10.4 214.4
VL2 L3.1 53.2 77.7 10.0 216.2
TABLE IV: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
Auxiliary Input Visual Textual $R@1\uparrow$ $R@5\uparrow$ $MeanR\downarrow$ $RSum\uparrow$
Captions Blip2 PG 51.2 76.2 11.2 212.4
Captions Blip2 L3.1 51.6 76.7 10.9 213.8
Captions VL2 L3.1 50.4 75.9 11.5 211.0
Tags VL2 L3.1 53.2 77.7 10.0 216.2
TABLE V: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
(a) Training and Inference.
(b) Inference.
Figure 5: Ablation on varying the number of tags across all inference strategies. $(N-M)$ on the x-axis indicates the number of training tags $N$ and inference tags $M$.
Note that we do not include methods that use extra input modalities during training or that rely on significantly stronger pre-training. For example, in [16, 7, 96, 93], models also train on audio data, so they are not directly comparable to MAC-VR, which only uses video and text. Moreover, VAST [96] includes subtitles (when available) in training as additional modality. This work is also pre-trained on a multi-modal video caption dataset. Similarly, CLIP-ViP [32] uses extra pre-training datasets to improve retrieval performance further. We also do not compare to T-MASS [97] as the official code has been retracted after a discovered data leakage.
Overall, MAC-VR achieves best overall performance when combined with DSL inference across 5 of the 6 datasets – only outperformed by [48] on DiDeMo. Results are consistent using QB inference for MSR-VTT (both splits) and without inference strategies on TGIF, Charades and YouCook2. On MSR-VTT (both splits), results are comparable to SOTA without an inference strategy.
Across 5 of the 6 datasets, MAC-VR shows the best performance for any inference strategy. However, on DiDeMo, its performance is lower than SOTA methods for all inference strategies, even though the DSL approach still gives good results. This difference may be influenced by different training parameters (i.e. $BS$ and $N_{v}$) used in other SOTA [15, 29, 80, 14, 48, 82, 83, 10]. To showcase this, we re-ran two SOTA methods [15, 82] using our training settings, and both performed worse than in their original experiments, i.e. -$\text{-}8.8$ for [15] and -$\text{-}15.3$ for [82], and worse than our MAC-VR, i.e. -$\text{-}0.6$ for [15] and -$\text{-}8.0$ for [82]. Another factor could be that we use the same foundation model settings (only 8 frames) to generate tags for every dataset. In the case of DiDeMo, we extract visual tags from the whole video and textual tags by concatenating individual captions. As a result, many frames may not match any labelled segments, adding irrelevant and detrimental information.
Figure 6: Qualitative results across all the datasets. Left Rank: The ranking results of our baseline without using auxiliary tags. Right Rank: The ranking results of MAC-VR, which incorporates extracted visual (VT) and textual (TT) tags to enhance retrieval. Blue and orange colours highlight meaningful tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
(c) w/o Modality-Specific Tags.
(d) w/ Modality-Specific Tags.
Figure 7: t-SNE plot of visual and textual concepts on MSR-VTT-9k and Charades with/without using auxiliary modality-specific tags.
IV-D Ablation Studies
We provide ablations on MAC-VR over: the number of tags; architecture design; choice of foundation models; and use of auxiliary captions vs. tags. All ablations are performed on the commonly used MSR-VTT-9k dataset. We provide full tables and additional ablations in the appx. (Secs. B and C respectively).
Num. of Tags in Training and Inference. In Fig. 5, we vary the number of visual/textual tags in training/inference – note we keep the same number of video/text tags. In Fig. 5(a), we adjust only the number of tags in training and use the same value during inference. MAC-VR without IS and with QB increases in performance until reaching the best performance with $12$ tags. However, for DSL, $RSum$ increases until the best performance of $RSum=215.8$ with $6$ tags, then the value begins to drop. When varying the number of tags only in inference, as shown in Fig. 5(b), we observe that $RSum$ increases rapidly and overcomes our baseline DiCoSA when using more than $2$ tags and getting the best performance with $RSum=216.2$ with $8$ tags using DSL. Best performance at inference is reported at $8$ tags with DSL and $12$ tags otherwise, showcasing that using more tags increases performance. The figure also shows that DSL consistently obtains the best performance compared to QB or no inference strategy. Accordingly, we use DSL in all remaining ablation studies.
Choice of Tag Modality In Tab. V, we ablate which modalities to extract tags from for MAC-VR. Using visual and textual tags individually gets results that are similar or slightly superior to our baseline. When combined, all metrics improve, outperforming our baseline.
Choice of Foundation Models. In Tab. V, we ablate the use of different foundation models to extract visual and textual tags from a video and its caption. We compare Video-LLaMA [38] against VideoLLaMA2 [37] and Llama2 [40] against Llama3.1 [95] to extract visual and textual tags respectively. Results show that all metrics improved using VideoLLaMA2 and Llama3.1 by $\Delta RSum=2.2$. Video-LLaMA and Llama2 tend to hallucinate tags more often, qualitative examples can be seen in Sec. D in appx.
Using Auxiliary Captions over Tags. In Tab. V, we show that tags are more informative compared to further captions extracted directly from the video or paraphrased from the caption. We use Blip2 [34]/VideoLLaMA2 [37] as video captioners and PEGASUS [98]/Llama3.1 [95] as paraphrasers, see appx. for details (Sec. E). Results show that using tags outperforms additional captions – including captions from the same models used to generate tags.
V Qualitative Results
Fig. 6 shows qualitative results on all datasets. We compare retrieved ranks of MAC-VR against our baseline without visual and textual tags. In general, both visual and textual tags add complementary information extracted from the video and the caption that directly help retrieval. For example, consider the query a class is being introduced to a digital reading device and its video top right of Fig. 6. Visual tags, student, technology, learning, and textual tags, initial setup, introduction, training, add additional information to describe the action shown in the video/described in the caption. We acknowledge there are cases where tags are harmful. For example, given the query a person is cooking on stage we can see that singing is one of the visual tags extracted. The VLM extracted this tag from the last frame where there is a person singing on stage as the scene changed. Additional results are shown in Sec. G in appx.
We align the visual and textual concepts with the corresponding auxiliary modality-specific concepts by introducing our Alignment Loss $\mathcal{L_{A}}$ (Sec. III-C). Fig. 7 shows the t-SNE of visual and textual concepts with/without the auxiliary tags on MSR-VTT and Charades ($1000$ samples), see Sec. F in the appx. for additional t-SNE plots. Our method better groups concepts into clusters over the baseline.
Discussion and Future Work. We find that modality-specific tags are certainly beneficial for video retrieval, but also discover that they can be harmful for specific queries. This could represent either correct tags that are redundant or incorrect tags due to errors in tag extraction. Foundation models, i.e. VLMs and LLMs, tend to hallucinate the content of the output meaning that the generated content might stray from factual reality or include fabricated information [99, 100]. Finally, we treat all the visual and textual tags with the same importance, it is possible that some generated words can be more or less discriminative than others based on uniqueness and/or relatedness to the video/caption. We leave this exploration for future work.
VI Conclusion
In this work, we introduce the notion of visual and textual tags extracted by foundation models from a video and its caption respectively and use them to boost the video retrieval performance. We propose MAC-VR (Modality Auxiliary Concepts for Video Retrieval), where we incorporate modality-specific auxiliary tags, projected into disentangled auxiliary concepts. We use a new Alignment Loss to better align each modality with its auxiliary concepts. We ablate our method to further show the benefit of using auxiliary modality-specific tags in video retrieval. Our results indicate, both qualitatively and by comparing to other approaches, that modality-specific tags help to decrease ambiguity in video retrieval on five video datasets.
Acknowledgements
This work used public datasets and was supported by EPSRC UMPIRE (EP/T004991/1) and EPSRC Program Grant Visual AI (EP/T028572/1).
Appendix A Tag Extraction
In Sec. III-B, we have described how we extracted tags from a video and its corresponding caption, here we provide additional information and the prompts used to extract tags for the video and text modalities. For all the datasets except DiDeMo, we use the video and its corresponding caption to generate visual and textual tags, respectively. For DiDeMo, due to it being used for video-paragraph retrieval, we first concatenate the captions from the video to form the associated paragraph. Then, we generate tags from the entire video and paragraph.
The prompt used as input of VideoLLaMA2 [37] to extract visual tags is: {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behavior or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action.
What are the top 10 general tags that capture the fundamental idea of this action? Give me a bullet list as output where each point is a general tag, and use one or two significant words per tag and do not give any explanation.
The prompt for Llama3.1 [95] to extract textual tags is: {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true] A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: You are a conversational AI agent. You typically extract general tags of an action.
A general tag of an action is a fundamental and overarching idea that encapsulates the essential principles, commonalities, or recurrent patterns within a specific behavior or activity, providing a higher-level understanding of the underlying themes and purpose associated with that action.
Given the following action: 1) {}
What are the top 10 general tags of the above action? Use one or two significant words per tag and do not give any explanation.
ASSISTANT:
Note that {} will be replaced with the caption. We did not provide any examples in the prompts (i.e. in context learning) as we found that this led to the foundation models hallucinating these examples in the output. Rather, we found that the foundation models were able to generate reasonable outputs for both video and text. We do not use any strategy to avoid the foundation models hallucinating as spot-checking the results found them to be clean enough for our purposes. The only post-processing strategy we adopted was to clean the output of the models in order to get the corresponding tags: we remove punctuation; stopwords; extracted tags that contain a noun and a verb to avoid the presence of complete sentences as tags; and tags larger than $3$ words. In Fig. 8 we show additional examples of tags for all the datasets.
$K$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
4 52.4 77.1 85.1 1 10.2 214.6
8 53.2 77.7 85.3 1 10.0 216.2
16 52.0 77.3 84.8 1 10.4 214.1
32 50.7 77.1 84.5 1 10.3 212.3
TABLE VI: Ablation on number of concepts $K$.
$\alpha_{1}$ $\alpha_{2}$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
0.5 1.0 52.5 76.5 85.6 1 11.0 214.6
1.0 1.0 53.2 77.7 85.3 1 10.0 216.2
2.0 1.0 52.6 77.2 85.8 1 10.5 215.6
5.0 1.0 52.0 76.8 85.3 1 11.1 214.1
1.0 0.0 52.1 77.3 85.9 1 12.9 215.3
1.0 0.5 53.1 76.7 85.5 1 10.0 215.3
1.0 1.0 53.2 77.7 85.3 1 10.0 216.2
1.0 2.0 53.0 77.5 85.4 1 10.1 215.9
1.0 5.0 52.1 76.7 85.6 1 10.9 214.4
1.0 10.0 52.5 77.5 85.4 1 10.5 215.4
TABLE VII: Ablation on loss parameters $\alpha_{1}$ and $\alpha_{2}$.
#Unique Tags Avg #Tags per item
Foundation Model Tags Train Test Train Test
Video-LLaMA [38] Visual 8,049 3,500 27,11 27,12
VideoLLaMA2 [37] Visual 63,383 12,118 27.69 27.83
Llama2 [40] Textual 162,571 5,058 14.96 15.20
Llama3.1 [95] Textual 320,351 8,326 27.17 26.52
TABLE VIII: Comparison of statistics of tags when using different foundation models on MSR-VTT-9k.
MSR-VTT-9k ($BS=128$, $N_{v}=12$) MSR-VTT-7k ($BS=128$, $N_{v}=12$) DiDeMo ($BS=64$, $N_{v}=50$)
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
DiCoSA* [3] - IJCAI’23 47.2 73.5 83.0 2 12.9 203.7 44.4 72.6 81.2 2 14.7 198.2 41.2 71.3 81.3 2 15.9 193.8
MAC-VR (ours) - - 48.8 74.4 83.7 2 12.3 206.9 45.3 71.9 81.8 2 15.0 199.0 43.4 72.5 82.3 2 16.9 198.4
DiCoSA* [3] QB IJCAI’23 48.0 74.6 84.3 2 12.9 206.9 44.7 73.3 81.3 2 14.3 199.3 43.7 73.2 81.7 2 16.8 198.6
MAC-VR (ours) QB - 49.3 75.9 83.5 2 12.3 208.7 45.6 72.7 82.1 2 16.0 200.4 45.5 74.8 82.3 2 16.2 202.6
DiCoSA* [3] DSL IJCAI’23 52.1 77.3 85.9 1 12.9 215.3 49.2 73.3 82.8 2 12.6 205.3 47.3 75.7 83.8 2 14.2 206.8
MAC-VR (ours) DSL - 53.2 77.7 85.3 1 10.0 216.2 49.8 74.6 84.3 2 12.7 208.7 50.2 76.2 84.2 1 15.1 209.6
TGIF ($BS=128$, $N_{v}=12$) Charades ($BS=128$, $N_{v}=12$) YouCook2 ($BS=128$, $N_{v}=12$)
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
DiCoSA* [3] - IJCAI’23 14.0 28.5 36.1 30 296.7 78.6 16.7 35.9 47.1 12 61.7 99.7 11.9 29.2 40.3 18 117.9 81.4
MAC-VR (ours) - - 14.2 28.8 36.7 28 294.4 79.7 17.8 40.0 51.4 10 55.0 109.2 12.3 30.7 42.3 16 106.7 85.3
DiCoSA* [3] QB IJCAI’23 14.1 28.4 36.4 30 313.6 78.9 18.8 39.5 49.0 11 60.8 107.3 12.9 31.5 42.3 16 108.7 86.7
MAC-VR (ours) QB - 14.8 29.7 37.5 28 289.9 82.0 19.4 40.5 52.2 10 55.4 112.1 12.9 31.4 43.4 16 107.1 87.7
DiCoSA* [3] DSL IJCAI’23 15.5 30.1 37.8 28 290.6 83.4 18.2 39.2 49.1 11 58.9 106.5 12.8 32.2 43.0 15 103.8 88.0
MAC-VR (ours) DSL - 16.2 31.2 39.0 25 284.3 86.4 19.5 42.1 54.4 8 52.4 116.0 13.1 32.1 43.4 15 99.3 88.6
TABLE IX: Comparison with baseline trained by using same training parameters of MAC-VR. our reproduced results with same training parameters. IS: Inference Strategy. $BS$: Batch Size. $N_{v}$: Number of Frames. The highest $RSum$ in each block is marked in bold.
Appendix B Full Tables of Experiments
Tab. IX details the results presented in Fig. 4 and reports Recall at $L=1,5,10$ ($R@L$), median rank $MR$, mean rank ($MeanR$) and $RSum$ for all the datasets when using all the inference strategies. Similarly, . X reports the full table of Tab. II. Tabs. XIIIXIII are the full tables of the ablation s. VV in the main paper.
MSR-VTT-9k MSR-VTT-7k DiDeMo
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
CLIP4Clip [77] - Neurocomp.’22 44.5 71.4 81.6 2 15.3 197.5 42.1 71.9 81.4 2 16.2 195.4 43.4 70.2 80.6 2 17.5 194.2
CenterCLIP [78] - SIGIR’22 44.2 71.6 82.1 2 15.1 197.9 43.7 71.3 80.8 2 16.9 195.8 - - - - - -
X-Pool [26] - CVPR’22 46.9 72.8 82.2 2 14.3 201.9 43.9 72.5 82.3 2 14.6 198.7 - - - - - -
TS2-Net [13] - ECCV’22 47.0 74.5 83.8 2 13.0 205.3 - - - - - - 41.8 71.6 82.0 2 14.8 195.4
EMCL-Net [15] - NeurIPS’22 46.8 73.1 83.1 2 - 203.0 - - - - - - 46.8 74.3 83.1 2 12.3 204.2
EMCL-Net [15] - NeurIPS’22 - - - - - - 45.2 71.4 81.2 2 15.0 197.8 44.1 71.7 80.3 2 15.1 197.8
VoP [27] - CVPR’23 44.6 69.9 80.3 2 16.3 194.8 42.7 68.2 79.3 2 15.9 190.2 46.4 71.9 81.5 2 13.6 199.8
MuMUR [79] - Inf. Retr. J.’23 46.4 72.6 82.2 2 13.9 201.2 44.8 72.0 82.5 2 - 199.3 44.4 74.3 83.1 2 - 201.8
PiDRo [29] - ICCV’23 48.2 74.9 83.3 2 12.6 206.4 - - - - - - 48.6 75.9 84.4 2 11.8 208.9
HBI [80] - CVPR’23 48.6 74.6 83.4 2 12.0 206.6 - - - - - - 46.9 74.9 82.7 2 12.1 204.5
DiffusionRet [14] - ICCV’23 49.0 75.2 82.7 2 12.1 206.9 - - - - - - 46.7 74.7 82.7 2 14.3 204.1
Prompt Switch [81] - ICCV’23 47.8 73.9 82.2 - 14.4 203.9 - - - - - - - - - - - -
Cap4Video [48] - CVPR’23 49.3 74.3 83.8 2 12.0 207.4 - - - - - - 52.0 79.4 87.5 1 10.5 218.9
UCoFiA [82] - ICCV’23 49.4 72.1 83.5 2 12.9 205.0 - - - - - - 46.5 74.8 84.4 2 13.4 205.7
UCoFiA [82] - ICCV’23 - - - - - - 45.2 69.6 79.4 2 15.7 194.2 42.1 69.2 79.1 2 16.3 190.4
PAU [83] - NeurIPS’23 48.5 72.7 82.5 2 14.0 203.7 - - - - - - 48.6 76.0 84.5 2 12.9 209.1
TABLE [10] - AAAI’23 47.1 74.3 82.9 2 13.4 204.3 - - - - - - 47.9 74.0 82.1 2 14.3 204.0
UATVR [33] - ICCV’23 47.5 73.9 83.5 2 12.3 204.9 - - - - - - 43.1 71.8 82.3 2 15.1 197.2
TeachCLIP [30] - CVPR’24 46.8 74.3 - - - - - - - - - - 43.7 71.2 - - - -
MV-Adapter [31] - CVPR’24 46.2 73.2 82.7 - - 202.1 - - - - - - 44.3 72.1 80.5 - - 196.9
BiC-Net [84] - TOMM’24 39.4 75.5 86.7 2 - 201.6 32.8 68.2 82.4 3 - 183.4 - - - - - -
RAP [85] - ACL’24 44.8 71.4 81.5 - 14.4 197.7 - - - - - - 42.6 70.4 79.6 - 18.0 192.6
MAC-VR (ours) - - 48.8 74.4 83.7 2 12.3 206.9 45.3 71.9 81.8 2 15.0 199.0 43.4 72.7 82.3 2 16.9 198.4
QB-Norm [17] QB CVPR’22 47.2 73.0 83.0 2 - 203.2 - - - - - - 43.3 71.4 80.8 2 - 195.5
DiCoSA [3] QB IJCAI’23 47.5 74.7 83.8 2 13.2 206.0 - - - - - - 45.7 74.6 83.5 2 11.7 203.8
DiffusionRet [14] QB ICCV’23 48.9 75.2 83.1 2 12.1 207.2 - - - - - - 48.9 75.5 83.3 2 14.1 207.7
MAC-VR (ours) QB - 49.3 75.9 83.5 2 12.3 208.7 45.6 72.7 82.1 2 16.0 200.4 45.5 74.8 82.3 2 16.2 202.6
EMCL-Net [15] DSL NeurIPS’22 51.6 78.1 85.3 1 - 215.0 - - - - - - - - - - - -
EMCL-Net [15] DSL NeurIPS’22 - - - - - - - - - - - - 47.6 73.5 82.8 2 11.9 203.9
TS2-Net [13] DSL ECCV’22 51.1 76.9 85.6 1 11.7 213.6 - - - - - - 47.4 74.1 82.4 2 12.9 203.9
TABLE [10] DSL AAAI’23 52.3 78.4 85.2 1 11.4 215.9 - - - - - - 49.1 75.6 82.9 2 14.8 207.6
UATVR [33] DSL ICCV’23 49.8 76.1 85.5 2 12.3 211.4 - - - - - - - - - - - -
RAP [85] DSL ACL’24 - - - - - - - - - - - - 47.1 74.1 82.4 - 13.9 203.6
MAC-VR (ours) DSL - 53.2 77.7 85.3 1 10.0 216.2 49.8 74.6 84.3 2 12.7 208.7 50.2 75.2 84.2 1 15.1 209.6
TGIF Charades YouCook2
Method IS Venue $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$ $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
RIVRL [86] - TCSVT’22 12.2 26.8 35.4 30 - 74.4 - - - - - - - - - - - -
X-Pool [26] - CVPR’22 - - - - - - 16.1 35.2 44.9 14 67.2 96.2 - - - - - -
AME-Net [87] - TOMM’22 - - - - - - - - - - - - 7.6 21.5 32.8 28 - 61.9
EMCL-Net [15] - NeurIPS’22 13.2 27.3 34.8 36 327.5 75.3 16.4 36.3 47.3 13 69.5 100.0 11.8 28.8 39.8 19 121.6 80.4
RoME [88] - CoRR’22 - - - - - - - - - - - - 6.3 16.9 25.2 53 - 48.4
SwAMP [89] - AISTATS’23 - - - - - - - - - - - - 9.4 24.9 35.3 22 - 69.6
MuMUR [79] - Inf. Retr. J.’23 - - - - - - 16.6 37.5 50.0 10 52.7 104.1 - - - - - -
UCoFiA [82] - ICCV’23 13.0 27.1 34.7 35 323.8 74.8 16.4 37.2 47.7 12 74.6 101.3 12.4 30.4 41.0 17 123.5 83.8
LTME [90] - ICASSP’24 10.6 24.1 32.2 - - 66.9 - - - - - - - - - - - -
BiC-Net [84] - TOMM’24 - - - - - - - - - - - - 8.7 23.9 33.5 31 - 66.1
MAC-VR (ours) - - 14.2 28.8 36.7 28 294.4 79.7 17.8 40.0 51.4 10 55.0 109.2 12.3 30.7 42.3 16 106.7 85.3
EMCL-Net [15] DSL NeurIPS’22 14.8 29.4 36.6 31 317.3 80.8 18.2 38.2 49.3 11 63.1 105.7 12.6 31.7 42.5 17 104.5 86.8
MAC-VR (ours) DSL - 16.2 31.2 39.0 25 284.3 86.4 19.5 42.1 54.4 8 52.4 116.0 13.1 32.1 43.4 15 99.3 88.6
TABLE X: Comparison with SOTA across all the datasets. --: unreported results. our reproduced results. IS: Inference Strategy. The highest $RSum$ in each block is marked in bold.
Method $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
baseline 52.1 77.3 85.9 1 12.9 215.3
+VT 52.2 77.3 85.7 1 10.4 215.2
+TT 52.0 77.6 86.0 1 10.4 215.6
+VT+TT 53.2 77.7 85.3 1 10.0 216.2
TABLE XI: Ablation on architecture design.
Foundation Models
VT TT $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
VL L2 52.0 77.5 84.9 1 10.4 214.4
VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
TABLE XII: Ablation on foundation models. VL: Video-LLaMA. VL2: VideoLLaMA2. L2: Llama2. L3.1: Llama3.1.
Auxiliary Input Visual Textual $R@1\uparrow$ $R@5\uparrow$ $R@10\uparrow$ $MR\downarrow$ $MeanR\downarrow$ $RSum\uparrow$
Captions Blip2 PG 51.2 76.2 85.0 1 11.2 212.4
Captions Blip2 L3.1 51.6 76.7 85.5 1 10.9 213.8
Captions VL2 L3.1 50.4 75.9 84.7 1 11.5 211.0
Tags VL2 L3.1 53.2 77.7 85.3 1 10.0 216.2
TABLE XIII: Ablation on auxiliary inputs. PG: PEGASUS. L3.1: Llama3.1. VL2: VideoLLaMA2.
Appendix C Additional Ablations
In this section, we ablate some additional design choices of MAC-VR. First, we ablate $K$ that controls the number of concepts/latent factors. We evaluate within the following range $K\in{4,8,16,32}$. In Tab. VIII, the performance improves reaching the best result with $K=8$ and then decreases. This shows that a small number of concepts limits the ability to leverage fine-grained information, whereas a larger value reduces the dimensionality of each concept limiting the discriminability of the concept itself.
In Eq. 5, we described our training loss and we introduced two weight parameters $\alpha_{1}$ and $\alpha_{2}$ that control the importance of the contrastive loss $L_{C}$ to align the modality concepts and the proposed Alignment Loss $L_{A}$ to align the modality with the auxiliary modality concepts. We consider different values of $\alpha_{1}\in\{0.5,1.0,2.0,5.0\}$ and $\alpha_{2}\in\{0.0,0.5,1.0,2.0,5.0,10.0\}$. We find that the best $RSum$ is when $\alpha_{1}=1.0$ and $\alpha_{2}=1.0$ with $RSum=216.2$. The performance drops when $\alpha_{1}>2.0$ or $\alpha_{2}>2.0$, as shown in Tab. VIII. These results show that both losses must contribute to the learning phase equally to better align visual and textual concepts and so improve the final retrieval performance.
Appendix D Differences between tags extracted from different foundations models
As explained in Sec. IV-D, we considered different foundation models to extract visual and textual tags. More precisely, we considered Video-LLaMA [38] and VideoLLaMA2 [37] to extract visual tags and Llama2 [40] and Llama3.1 [95] to extract textual tags. We used the same parameters and same prompts to extract the tags by using all the considered foundations models.
Tab. VIII shows the statistics of extracted tags for MSR-VTT-9k. The total number of unique extracted visual tags by using Video-LLaMA [38] is much smaller than the total number obtained when using VideoLLaMA2 [37]. The same conclusion is reached for the total number of unique extracted textual tags by Llama2 [40] and Llama3.1 [95] respectively. This shows the lesser ability of older foundation models to generate tags compared with more recent versions. In particular, this is more evident when we focused on the visual tags, where not only the total number of unique tags is smaller but also the average number of tags per pairs is the same, meaning that many tags are shared among pairs and so there are less unique tags able to distinguish all the video/caption pairs.
Fig. 9 shows a qualitative comparison of the extracted tags by using different foundation models. It is evident that Video-LLaMA and Llama2 tend to hallucinate tags that are not relevant with what shown in the video and described in the caption. Moreover, the textual tags extracted by using Llama2 are very often words that already appear in the caption.
Figure 9: Comparison of extracted visual and textual tags on the MSR-VTT dataset when using different foundation models. Blue and orange colours highlight meaningful tags.
For example, given the captions a cartoon character runs around inside of a video game and its corresponding video, we can see that Video-LLaMA and Llama2 hallucinate some visual tags such as snowboarding, desert, skiing, bird, climbing—definitely irrelevant to what appear in the video—and textual tags such as video game, running, character, game character that are already words that appear in the caption, therefore they do not add any additional information to better retrieve the correct video. On the contrary VideoLLaMA2 and Llama3.1 tend to extract tags that add additional information to the video and text. See Fig. 9 for more examples on all the considered datasets.
Appendix E How to Generate Auxiliary Captions.
In Sec. IV-D, we ablate the use of auxiliary captions instead of using tags. We generate these additional captions by extracting them directly from the video and paraphrasing the original caption. We consider different approaches to extract captions from video and text.
Visual Captions. We consider two different approaches to generate new captions from a video: Blip2 [34] and VideoLLaMA2 [37]. Following the same approach proposed in [101], we generate new captions by extracting the middle frame of each video and use Blip2 to generate a new caption. We use the same parameters of VideoLLaMA2 to extract captions as we did to extract visual tags in MAC-VR, as described in Sec. IV-B. However, we used a general prompt to ask VideoLLaMA2 to generate new captions. {mdframed}[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
You are a conversational AI agent. You typically look at a video and generate a new caption for a video. Generate 10 new captions. Give me a bullet list as output.
Textual Captions. We consider the paraphraser PEGASUS [98] and Llama3.1 [95] to paraphrase the original caption. PEGASUS [98] is a standard Transformer-based encoder-decoder method pre-trained on a massive text corpora with a novel pre-training objective called Gap Sentence Generation (GSG). Instead of using traditional language modeling, PEGASUS removes important sentences from a document (gap-sentences) and asks the model to predict these missing sentences. After the pre-training stage, PEGASUS is fine-tuned on specific summarization datasets to improve its performance on downstream tasks. The model becomes highly effective at generating concise and accurate summaries by leveraging its pre-training knowledge. We extract new captions from a caption by Llama3.1 [95] by giving as input a general prompt as we did to extract tags:
{mdframed}
[backgroundcolor=blue!10, linecolor=red, linewidth=2pt, roundcorner=10pt, shadow=true]
A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. USER: You are a conversational AI agent. You typically paraphrase sentences by using different words but keeping the same meaning.
Given the following sentence: 1) {}
Generate 10 different sentences that are a paraphrased version of the original sentence. Give me a bullet list as output.
ASSISTANT:
We randomly pick an extracted visual and textual caption as auxiliary inputs in MAC-VR during training, noting a similar length of input with our ‘sentence of tags’. In inference, we always pick the first caption in the set of the extracted ones. Some examples of the extracted captions with all the considered methods are shown in Fig. 10.
Figure 10: Examples of extracted captions on the MSR-VTT dataset.
Appendix F Additional t-SNE plot.
In Fig. 12, we show the t-SNE plot of MAC-VR on the MSR-VTT test set when using modality-specific tags with/without our Alignment Loss $\mathcal{L}_{A}$. It is clear that the use of $\mathcal{L}_{A}$ helps to better distinguish the different concepts and have better clusters in the t-SNE plot. In Fig. 13 and Fig. 14, we show the t-SNE plot of visual and textual concepts without auxiliary modality-specific tags and when using only visual tags (Fig. 13) and textual tags (Fig. 14) with our Alignment Loss $\mathcal{L}_{A}$. From this, both tags used individually help to better align the visual and textual concepts. In particular, the visual tags help to better align the visual and textual concepts compared to the textual tags. A possible explanation for this is that the tags extracted from videos share the same modality as captions, which facilitates better alignment between visual and textual concepts. We leave this conclusion as possible inspiration for future works in this field.
Fig. 15 to 17 show the t-SNE plot of MAC-VR of visual and textual concepts with/without auxiliary modality-specific tags across the other datasets. We sample $1,000$ pairs for TGIF and YouCook2. We can see that the use of auxiliary modality-specific tags help to better distinguish the different concepts and have better clusters in the t-SNE plot.
Appendix G Qualitative Results
In Fig. 11, we show additional qualitative results across all the datasets. It is evident that tags add additional information that is not described in the text. For example, the video associated to the caption a guy is in the water with his surfboard has visual (i.e. water activity, ocean, waves, surfing, outdoor activity) and textual (i.e. surfing, water, ocean, riding) tags that add additional information and boost the rank position by $3$. Similarly, in the YouCook2 example, the video is associated to a general mash the beans, tags help to add additional information, indeed visual tags such as boiling, homemade soup, boiling liquid, thickening sauce and textual tags such as such as crushing action, bean processing, pureeing, ingredient mashing add semantic information that definitely help the model in the retrieval process. Rarely, the extracted tags were not semantically relevant with what shown in the video or described in the caption. For instance, some visual and textual tags of the video associated to the caption the man blows out the candles. the girl puts the cake down. the camera pans to a man blowing out the candles on his cake. man blows candles out a plate is placed on the table are marriage partnership, love, love commitment as visual tags and class as textual tags. In other cases, visual and textual tags are too general, even though they are relevant to the video and caption, and so they do not add any additional semantic information. For example the video associated to the caption place wasabi next to the tuna are food, cooking, preparation techniques as visual tags and ingredients, serving, serving food. Even though these tags are relevant to the video and caption, they are too general and do not add any additional semantic knowledge.
Figure 11: Additional qualitative results across datasets. Blue and orange colours highlight some meaningful tags.
(a) w/ Modality-Specific Tags, w/o $\mathcal{L}_{A}$.
(b) w/ Modality-Specific Tags, w/ $\mathcal{L}_{A}$.
Figure 12: t-SNE plot of visual and textual concepts on the MSR-VTT-9k with/without the Alignment Loss $\mathcal{L}_{A}$ with using both visual and textual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Visual Tags, w/ $\mathcal{L}_{A}$.
Figure 13: t-SNE plot of visual and textual concepts on the MSR-VTT-9k without using auxiliary modality-specific tags and with using only visual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Textual Tags, w/ $\mathcal{L}_{A}$.
Figure 14: t-SNE plot of visual and textual concepts on the MSR-VTT-9k without using auxiliary modality-specific tags and with using only textual tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
Figure 15: t-SNE plot of visual and textual concepts on the DiDeMo with/without using auxiliary modality-specific tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
Figure 16: t-SNE plot of visual and textual concepts on the TGIF with/without using auxiliary modality-specific tags.
(a) w/o Modality-Specific Tags.
(b) w/ Modality-Specific Tags.
Figure 17: t-SNE plot of visual and textual concepts on the YouCook2 with/without using auxiliary modality-specific tags.